# Never-Ending Behavior-Cloning Agent for Robotic Manipulation

### v2

@@ Abstract
Relying on multi-modal observations, embodied robots could perform multiple robotic manipulation tasks in unstructured real-world environments. However, most language-conditioned behavior-cloning agents still face existing long-standing challenges, i.e., 3D scene representation and human-level task learning, when adapting into new sequential tasks in practical scenarios. We here investigate these above challenges with NBAgent in embodied robots, a pioneering language-conditioned Never-ending Behavior-cloning Agent. It can continually learn observation knowledge of novel 3D scene semantics and robot manipulation skills from skill-shared and skill-specific attributes, respectively. Specifically, we propose a skill-shared semantic rendering module and a skill-shared representation distillation module to effectively learn 3D scene semantics from skill-shared attribute, further tackling 3D scene representation overlooking. Meanwhile, we establish a skill-specific evolving planner to perform manipulation knowledge decoupling, which can continually embed novel skill-specific knowledge like human from latent and low-rank space. Finally, we design a never-ending embodied robot manipulation benchmark, and expensive experiments demonstrate the significant performance of our method.
*Corresponding Author.
1 Introduction
@@
Robot learning has attracted growing interests in integrating machine learning with robot control system to solve various robotic tasks, such as manipulation [39], navigation [36], mapping and localization [27]. In recent decades, behavior-cloning agents [17, 53] have demonstrated notable success in effective training with only a few demonstrations and direct implementation on real robots. Benefiting from vision-language observations, current multi-task behavior-cloning methods focus on utilizing multi-modal data to efficiently execute complex manipulation tasks with visual observations. For instance, PerAct [39] utilizes a PerceiverIO Transformer [21] to encode language goals and RGB-D voxel observations, subsequently generating discretized robotic actions.
@@
However, most existing language-conditioned behavior-cloning methods [21, 17, 26] assume that the skills mastered by a robot remain unchanged over time in real-world 2D scenarios, and focus on training policy model on a fixed set of manipulation tasks. They are resource-constrained when undertaking new skills and handling novel objects with intricate structures [44]. To address this scenario, the embodied robots—like human—should be capable of learning novel challenging manipulation skills in a continual learning manner. For instance, a home robot in the open-ended world is expected to consecutively learn various novel manipulation skills with demonstrations of behavior, and timely meet the evolving needs of its owners. Generally, a trivial approach for this scenario involves retraining the robot on all past and novel data, which leads to a large computational burden and high cost of memory storage, thereby is limited in real world unstructured 3D applications.
@@
To address the aforementioned scenarios, we consider a practical challenging robot learning problem, i.e., Never-ending Behavior-cloning Robot Learning (NBRL), where an agent can perform continual learning on successive behavior-cloning manipulation tasks and counteract catastrophic forgetting from learned skills. Different from learning category-wise knowledge focused by most existing methods [33, 44, 11], we here take the attempt to rethink some attributes about skill-wise knowledge:
@@
•  Skill-Shared Attribute indicates that different complex skills possess shared knowledge, such as similar manipulated objects and 3D scenes understanding and so on. We primarily investigate the consistency of skill-shared knowledge across various skills on semantics of 3D scenes, which plays a key role in addressing skill forgetting and achieving comprehensive understanding on 3D scenes.
@@
•  Skill-Specific Attribute originates from various factors such as distinct manipulation sequences, object recognition, motion primitives, and other distinct elements within each unique skill. Focusing the embodied robot to learn skill-specific knowledge can effectively perform novel skills learning and tackle forgetting on past learned manipulation skills or tasks.
@@
To tackle the above-mentioned challenges, we propose a pioneering language-conditioned Never-ending Behavior-cloning Agent (i.e., NBAgent), which can continually acquire skill-wise knowledge from both skill-shared and skill-specific attributes. To the best of our knowledge, this is an earlier attempt to explore continual learning for multi-modal behavior-cloning robotic manipulation. To be specific, we design a skill-shared semantic rendering module (SSR) and a skill-shared representation distillation module (SRD) to transfer skill-shared knowledge on semantics of 3D scenes. Supervised by Neural Radiance Fields (NeRFs) and a vision foundation model, SSR can transfer skill-shared semantic from 2D space into 3D space across novel and old skills; SRD can effectively distill skill-shared knowledge between old and current models to well align voxel representation. Additionally, we propose a skill-specific evolving planner (SEP) to decouple the skill-wise knowledge in the latent and low-rank space, and focus on skill-specific knowledge learning to alleviate manipulation skill forgetting. Several major contributions of our work are as follows:
@@
•  We develop a skill-shared semantic rendering module and a skill-shared representation distillation module to capture the 3D semantic knowledge, which can transfer skill-shared semantic of 3D scenes to overcome 3D reasoning overlooking in continual learning.
@@
•  We design a skill-specific evolving planner to perform skill-specific knowledge learning, which can decouple the skill-wise knowledge into latent and low-rank space, and continually embed novel skill-specific knowledge. Comprehensive experiments conducted on our proposed NBRL benchmark for home robotic manipulation demonstrate the effectiveness and robustness of our method.
2 Related Work
Robotic Manipulation. Recent works [6, 8, 20, 37, 14, 5, 54] have resulted in substantial advancements in the accomplishment of intricate tasks. VIMA [24] proposes a novel multi-modal prompting scheme, which transforms a wide range of robotic manipulation tasks into a sequence modeling problem. Compared with directly using images as the manipulation input [16, 38], voxelizing 3D point clouds as a 3D representation [39, 53, 17, 23] can accomplish more complex tasks. PerAct [39] enables agent to perform better in robotic manipulation by voxelizing RGB-D images and discretizing output actions. GNFactor [53] can generalize neural radiance field rendering through joint training, improving the model’s generalization ability.
Continual Learning. Continual learning provides the foundation for the adaptive development of AI systems [46, 10]. The main approaches of continual learning can be categorized into three directions: Parameter regularization-based methods [33, 30, 9, 12, 11] balance the old and new tasks by adding more explicit regularization terms. Architecture-based methods [25, 48, 45, 43] construct network parameters for different tasks. Replay-based methods include empirical replays [4, 33, 41, 42] and generative replays [29, 49, 40]. Some works focus on the improvement of robots by continual learning [3, 15, 18, 2, 1]. LOTUS [44] stores a few human demos of novel tasks into the growing skill library, enabling lifelong learning ability for robots. However, these methods are build on 2D vision observations, and cannot be applied to perform language-conditioned behaviour-cloning manipulation with multi-modal data.
3 Methodology
3.1 Problem Definition and Overview
@@
Figure 1: Overview of the proposed NBAgent. It consists of a skill-shared semantic rendering module and a skill-shared representation distillation loss $\mathcal{L}_{\mathrm{SRD}}$ to transfer skill-shared knowledge on semantics of 3D scenes, and a skill-specific evolving planner to learn skill-specific knowledge.
@@
3.2 Skill-Shared Semantic Rendering Module
@@
[…]
@@
$\mathcal{L}_{\mathrm{C}}=\underset{\mathbf{r}\in\mathcal{R}}{\sum}\|\mathbf{C}(\mathbf{r},\mathbf{v}_{s})-\mathbf{Y}_{c}(\mathbf{r})\|^{2}_{2}+\beta\underset{\mathbf{r}\in\mathcal{R}}{\sum}\|\mathbf{C}(\mathbf{r},\mathbf{v}_{s})-\hat{\mathbf{C}}(\mathbf{r},\mathbf{v}_{s})\|^{2}_{2}\cdot\mathbb{I}_{\mathbf{v}_{s}\notin\mathcal{T}^{m}},$ (3)
@@
$\mathbf{Y}_{c,t}(\mathbf{r}):=\sqrt{\alpha_{t}}\mathcal{E}_{v}(\mathbf{Y}_{c}(\mathbf{r}))+\sqrt{1-\alpha_{t}}\epsilon,~{}~{}\hat{\mathbf{F}}_{s}(\mathbf{r},\mathbf{l}_{p})=\Theta_{u}(\mathbf{Y}_{c,t}(\mathbf{r}),\mathcal{E}_{c}(\mathbf{l}_{p})),$ (4)
where $\mathcal{E}_{v}$ is a VAE encoder to encode image $\mathbf{Y}_{c}$ from pixel space to latent semantic space. $t$ represents the diffusion process step, $\epsilon\sim\mathcal{N}(0,1)$ and $\alpha_{t}$ is designed to control the noise schedule. $\mathbf{l}_{p}$ denotes the language prompt modified from task description $\mathbf{l}$. To perform skill-shared semantic transfer, we align the rendered semantic feature $\mathbf{F}_{s}$ and diffusion feature $\hat{\mathbf{F}}_{s}$, and obtain the major objective of our SSR module as follows:
$\mathcal{L}_{\mathrm{SSR}}=\mathcal{L}_{\mathrm{C}}+\lambda_{1}\mathcal{L}_{\mathrm{S}},~{}~{}\mathcal{L}_{\mathrm{S}}=\underset{\mathbf{r}\in\mathcal{R}}{\sum}\|\mathbf{F}_{s}(\mathbf{r},\mathbf{v}_{s})-\hat{\mathbf{F}}_{s}(\mathbf{r},\mathbf{l}_{p})\|^{2}_{2}.$ (5)
Algorithm 1 Optimization Pipeline of NBAgent.
0:  Continual skill learning tasks $\{\mathcal{T}^{m}\}_{m=1}^{M}$ with datasets $\mathcal{T}^{m}=\{\mathcal{D}^{m}\}_{i=1}^{N^{d}}$; Initialized perceiver model: $\Theta^{0}_{p}$; Initialized NeRF model: $\Theta^{0}_{n}$; Pre-trained diffusion model: $\Theta_{u}$; Pre-trained CLIP language encoder: $\mathcal{E}_{c}$; Initialized memory buff: $\mathcal{M}=\emptyset$; Iterations: $\{\mathcal{I}^{m}\}_{m=1}^{M}$.
1:  #While observing a new task $\mathcal{T}^{m}$:
2:  for $z=1,2,\cdots,\mathcal{I}^{m}$ do
3:     Randomly select keyframe $\mathbf{k}^{m}_{j}$ from $\{\mathcal{D}_{i}^{m}\}_{i=1}^{N^{d}}\cup\mathcal{M}$;
4:     Obtain $\mathrm{v}^{m},\mathrm{v}^{m-1},\mathrm{l}_{s},\mathrm{l}_{x}$ utilizing $\mathcal{E}^{m}_{v}$, $\mathcal{E}^{m-1}_{v}$and $\mathcal{E}_{c}$;
5:     Compute $\mathcal{L}_{\mathrm{SSR}}$ by SSR ($\mathrm{v}^{m},\mathrm{v}^{m-1},\mathrm{l}_{x},\Theta_{u}$) using Eq. (1);
6:     $\mathbf{S}[h,:],\mathbf{W}[h,:]\xleftarrow{}$ SEP ($\mathbf{l}_{s}$);
7:     Compute $\mathcal{L}_{\mathrm{CE}},\mathcal{L}_{\mathrm{SRD}}$ utilizing $\mathbf{S}[h,:]$, $\mathbf{W}[h,:]$, $\Theta^{m}_{p},\Theta^{m-1}_{p}$;
8:     Update $\Theta^{m}_{p}$ by Eq. (11);
9:  end for
10:  Store few samples from $\{\mathcal{D}_{i}^{m}\}_{i=1}^{N^{d}}$ in $\mathcal{M}$;
11:  Return: $\Theta^{m}_{p},\mathcal{M}$.
3.3 Skill-Shared Representation Distillation Module

### v3

@@ Abstract
Relying on multi-modal observations, embodied robots (e.g., humanoid robots) could perform multiple robotic manipulation tasks in unstructured real-world environments. However, most language-conditioned behavior-cloning agents in robots still face existing long-standing challenges, i.e., 3D scene representation and human-level task learning, when adapting into a series of new tasks in practical scenarios. We here investigate these above challenges with NBAgent in embodied robots, a pioneering language-conditioned Never-ending Behavior-cloning Agent, which can continually learn observation knowledge of novel 3D scene semantics and robot manipulation skills from skill-shared and skill-specific attributes, respectively. Specifically, we propose a skill-shared semantic rendering module and a skill-shared representation distillation module to effectively learn 3D scene semantics from skill-shared attribute, further tackling 3D scene representation overlooking. Meanwhile, we establish a skill-specific evolving planner to perform manipulation knowledge decoupling, which can continually embed novel skill-specific knowledge like human from latent and low-rank space. Finally, we design a never-ending embodied robot manipulation benchmark, and expensive experiments demonstrate the significant performance of our method.
I Introduction
@@
Robot learning has attracted growing interest in integrating machine learning with a robot control system to solve various robotic tasks, such as manipulation [53], navigation [50], mapping and localization [31]. In recent decades, behavior-cloning agents [19, 70] have demonstrated notable success in effective training with only a few demonstrations and direct implementation in real robots. Benefiting from vision-language observations, current multi-task behavior-cloning methods focus on utilizing multi-modal data to efficiently execute complex manipulation tasks with visual observations. For instance, PerAct [53] utilizes a PerceiverIO Transformer [23] to encode language goals and RGB-D voxel observations, subsequently generating discretized robotic actions. 3D Diffuser Actor [30] introduces a novel 3D denoising transformer that integrates 3D visual scene representations, language instructions, and proprioception to predict noise in 3D robot pose trajectories.
Figure 1: Demonstration illustration of our proposed never-ending behavior-cloning robot learning. As illustrated in (a), behavior-cloning robot learning primarily focuses on initially training on a fixed dataset, subsequently relying on the generalization capability to execute tasks in unseen environments, where a pre-trained CLIP model [45] serves as the the language encoder to process the language instruction. As depicted in (b), the never-ending behavior-cloning framework enables robotic systems to progressively acquire novel manipulation skills in a continual learning manner, thereby demonstrating enhanced adaptability and generalization capabilities when confronted with unseen and challenging tasks.
@@
However, most existing language-conditioned behavior-cloning methods [23, 19, 30] assume that the skills mastered by a robot remain unchanged over time in real-world 2D scenarios, and focus on training policy model on a fixed set of manipulation tasks. As shown in Fig. 1(a), they are generalization-constrained when undertaking new skills and handling novel objects with intricate structures [60]. To address this scenario, the embodied robots—like human—should be capable of learning novel challenging manipulation skills in a continual learning manner. For instance, a home robot in the open-ended world is expected to consecutively learn various novel manipulation skills with demonstrations of behavior, and timely meet the evolving needs of its owners. Generally, a trivial approach for this scenario involves retraining the robot on all past and novel data, which leads to a large computational burden and high cost of memory storage, thereby is limited in real-world unstructured 3D applications.
@@
To tackle the aforementioned challenges in real-world robotic applications, we formulate and introduce a practically significant problem framework, i.e., Never-ending Behavior-cloning Robot Learning (NBRL). As conceptually depicted in Fig. 1(b), the proposed NBRL framework empowers embodied agents to engage in never-ending skill acquisition through continual learning, while effectively mitigating catastrophic forgetting on learned skills. Different from learning category-wise knowledge focused by most existing methods [46, 60, 12], skill-wise knowledge encompasses object understanding, scene understanding, action semantics, language grounding, and operation sequences. We here take the attempt to rethink some attributes about skill-wise knowledge:
@@
•  Skill-Shared Attribute indicates that different complex skills share common knowledge in object understanding, scene understanding, action semantics, language grounding, and operation sequences. For example, the action semantics of pouring and the scene understanding required in both pouring water into a cup and pouring juice into a bowl illustrate this shared knowledge. We primarily investigate the consistency of skill-shared knowledge across various skills on semantics of 3D scenes, which plays a key role in addressing skill forgetting and achieving comprehensive understanding on 3D scenes.
@@
•  Skill-Specific Attribute denotes the unique knowledge components associated with different skills, including object understanding, scene understanding, action semantics, language grounding, and operation sequences, e.g., distinguishing the object understanding of a cup versus a bowl, or task-specific variations in operation sequences. Focusing the embodied robot to learn skill-specific knowledge can effectively perform novel skills learning and tackle forgetting on past learned manipulation skills or tasks.
@@
To tackle the above-mentioned challenges, we propose a pioneering language-conditioned Never-ending Behavior-cloning Agent (i.e., NBAgent), which can continually acquire skill-wise knowledge from both skill-shared and skill-specific attributes. To the best of our knowledge, this is an earlier attempt to explore continual learning for multi-modal behavior-cloning robotic manipulation. To be specific, we design a skill-shared semantic rendering module (SSR) and a skill-shared representation distillation module (SRD) to transfer skill-shared knowledge on semantics of 3D scenes. Supervised by Neural Radiance Fields (NeRFs) and a vision foundation model, SSR can transfer skill-shared semantic from 2D space into 3D space across novel and old skills. Meanwhile, SRD can effectively distill skill-shared knowledge between old and current models to well align voxel representation. Additionally, we propose a skill-specific evolving planner (SEP) to decouple the skill-wise knowledge in a latent and low-rank space, and focus on skill-specific knowledge learning to alleviate manipulation skill forgetting. Extensive experiments on robotic manipulation tasks validate that our proposed NBAgent model achieves state-of-the-art performance in NBRL problem.
Several major contributions of our work are as follows:
@@
  • We develop a skill-shared semantic rendering module and a skill-shared representation distillation module to capture the 3D semantic knowledge, which can transfer skill-shared semantic of 3D scenes to overcome 3D reasoning overlooking in continual skill learning.
@@
  • We design a skill-specific evolving planner to perform skill-specific knowledge learning, which can decouple the skill-wise knowledge into a latent and low-rank space, and continually embed novel skill-specific knowledge. Comprehensive experiments conducted on our proposed NBRL benchmark for home robotic manipulation demonstrate the effectiveness and robustness of our method.
Figure 2: Overview of the proposed NBAgent. It consists of a skill-shared semantic rendering module and a skill-shared representation distillation loss $\mathcal{L}_{\mathrm{SRD}}$ to transfer skill-shared knowledge on semantics of 3D scenes and overcome 3D reasoning overlooking in continual learning, and a skill-specific evolving planner to learn skill-specific knowledge, addressing catastrophic forgetting on learned skills.
II Related Work
Robotic Manipulation. Recent works [6, 36, 10, 22, 51, 16, 5, 72] have resulted in substantial advancements in robotic manipulation utilizing mult-imodal observation. Generalizable observation representation and language understanding are the key factors in achieving effective task reasoning and robust policy execution. VIMA [27] proposes a novel multi-modal prompting scheme, which transforms a wide range of robotic manipulation tasks into a sequence modeling problem. Octo[57] enhances robotic manipulation with a transformer-based policy, supporting versatile task definitions and efficient fine-tuning. GPT-4V(ision)[59] enables robots to perform tasks from human demonstrations by generating executable programs from video inputs. Compared with directly using images as the manipulation input [18, 52], voxelizing 3D point clouds as a 3D representation [53, 70, 19, 25] can accomplish more complex tasks. PPM6D [69] performs differentiable 6-D pose estimation for robotic grasping by cross-fusing RGB–point-cloud features with attention-guided point-pair matching, while [39] develops a consensus-driven distributed off-policy actor–critic algorithm with Lyapunov-based convergence guarantees and validates it on multi-UR5 robotic arm control. GNFactor [70] can generalize neural radiance field rendering to enhance 3D representation through joint training, improving the model’s generalization ability. ManiGaussian[40] enhances robotic manipulation using dynamic Gaussian Splatting to model spatiotemporal dynamics and predict optimal actions via future scene reconstruction. 3D Diffuser Actor[30] combines 3D scene representations with diffusion models to achieve SOTA performance in robotic manipulation.
Continual Learning. Continual learning provides the foundation for the adaptive development of AI systems [62, 71, 13]. The main approaches of continual learning can be categorized into three directions: Parameter regularization-based methods [46, 35, 11, 14, 12] balance the old and new tasks by adding more explicit regularization terms. C-LoRA[54] reduces catastrophic forgetting in diffusion models via self-regularized low-rank adaptation. RTRA[44] accelerates continual learning training with natural gradient descent, improving efficiency while maintaining effective performance. Architecture-based methods [29, 64, 61, 58] construct network parameters for different tasks. AdaB2N [41] enhances continual learning through adaptive Batch Normalization, improving stability and performance. InfLoRA [38] enhances continual learning by injecting a small number of parameters to reparameterize pre-trained weights and designing a subspace to eliminate interference between tasks. Replay-based methods include empirical replays [4, 34, 56, 37] and generative replays [33, 65, 55]. SDDR[28] improves class-incremental learning by generating synthetic samples with Stable Diffusion for distillation and replay. Some works focus on the improvement of robots by continual learning [3, 17, 20, 2, 1]. [26] proposes an end-to-end multi-object grasping policy for a musculoskeletal arm with object-preference experience replay to mitigate forgetting, while LEGION [42] introduces a Bayesian nonparametric knowledge-space framework augmented with language embeddings to accumulate and reuse lifelong RL skills for long-horizon robotic tasks. LOTUS [60] stores a few human demos of novel tasks into the growing skill library, enabling lifelong learning ability for robots. M2Distill [49] is a multimodal distillation framework for lifelong imitation learning that enforces cross-step consistency of vision–language–action latent representations and constrains consecutive Gaussian Mixture Model policies. However, these methods predominantly rely on 2D visual observations, and thus lack explicit 3D representations, which fundamentally limits spatial reasoning under occlusion changes and hinders robust language–action grounding for scalable language-conditioned behavior cloning. Moreover, many existing methods treat skill-wise knowledge as category-wise knowledge, overlooking its intrinsic decomposition into skill-shared and skill-specific components, which limits transfer across skills and exacerbates forgetting.
III Methodology
III-A Problem Definition and Overview
@@
III-B Skill-Shared Semantic Rendering Module
@@
where $\mathcal{F}_{\Theta^{m}_{n}}$ denotes the neural rendering function of NeRF model $\Theta^{m}_{n}$ in the $m$-th continual skill learning task. The 3D voxel feature $\mathbf{v}_{s}$ is obtained by a grid sample method based on trilinear interpolation from the 3D voxel observation $\mathbf{v}$. $\mathbf{x}$ and $\sigma$ are the 3D input point and differential density, and $\mathbf{d}$ represents unit viewing direction. The camera ray $\mathbf{r}$ can be obtained by: $\mathbf{r}=\mathbf{o}+t\mathbf{d}$, where $\mathbf{o}$ indicates the camera origin. By adding field-wise branches, $\Theta^{m}_{n}$ performs the same neural rendering function $\mathcal{F}_{\Theta^{m}_{n}}$ to estimate RGB color $\mathbf{c}$ and semantic feature $\mathbf{s}$. Therefore, the same accumulated transmittance $T(t)$ is shared to predict the two different fields and is defined as $T(t)=\mathrm{exp}(-\int^{t}_{t_{n}}\sigma(s)ds)$. In light of this, a RGB image $\mathbf{C}$ and 2D semantic map $\mathbf{M}$ can be rendered as :
$\mathbf{C}(\mathbf{r},\mathbf{v}_{s})=\int^{t_{f}}_{t_{n}}T(t)\sigma(\mathbf{r},\mathbf{v}_{s})\mathbf{c}(\mathbf{r},\mathbf{d},\mathbf{v}_{s})dt,$ (2)
$\mathbf{M}(\mathbf{r},\mathbf{v}_{s})=\int^{t_{f}}_{t_{n}}T(t)\sigma(\mathbf{r},\mathbf{v}_{s})\mathbf{s}(\mathbf{r},\mathbf{d},\mathbf{v}_{s})dt.$ (3)
Initialize: Continual skill learning tasks $\{\mathcal{T}^{m}\}_{m=1}^{M}$ with datasets $\mathcal{T}^{m}=\{\mathcal{D}^{m}\}_{i=1}^{N^{d}}$; perceiver model: $\Theta^{0}_{p}$; NeRF model: $\Theta^{0}_{n}$; Pre-trained diffusion model: $\Theta_{u}$; Pre-trained CLIP language encoder: $\mathcal{E}_{c}$; Memory buff: $\mathcal{M}=\emptyset$; Iterations: $\{\mathcal{I}^{m}\}_{m=1}^{M}$.
$\triangleright$ While observing a new task $\mathcal{T}^{m}$:
for $z=1,2,\cdots,\mathcal{I}^{m}$ do
  Randomly select keyframe $\mathbf{k}^{m}_{j}$ from $\{\mathcal{D}_{i}^{m}\}_{i=1}^{N^{d}}\cup\mathcal{M}$;
  Obtain $\mathrm{v}^{m},\mathrm{v}^{m-1},\mathrm{l}_{s},\mathrm{l}_{x}$ utilizing $\mathcal{E}^{m}_{v}$, $\mathcal{E}^{m-1}_{v}$and $\mathcal{E}_{c}$;
  Compute $\mathcal{L}_{\mathrm{SSR}}$ by SSR ($\mathrm{v}^{m},\mathrm{v}^{m-1},\mathrm{l}_{x},\Theta_{u}$) in Eq. (7);
 $\mathbf{S}[h,:],\mathbf{W}_{r}[h,:]\xleftarrow{}$ SEP ($\mathbf{l}_{s}$);
  Compute $\mathcal{L}_{\mathrm{CE}},\mathcal{L}_{\mathrm{SRD}}$ utilizing $\mathbf{S}[h,:]$, $\mathbf{W}_{r}[h,:]$, $\Theta^{m}_{p},\Theta^{m-1}_{p}$;
  Update $\Theta^{m}_{p}$ by Eq. (14);
Store few samples from $\{\mathcal{D}_{i}^{m}\}_{i=1}^{N^{d}}$ in $\mathcal{M}$;
Return: $\Theta^{m}_{p},\mathcal{M}$.
Algorithm 1 Optimization Pipeline of NBAgent.
@@
$\mathcal{L}_{\mathrm{C}}$ $=\underset{\mathbf{r}\in\mathcal{R}}{\sum}\|\mathbf{C}(\mathbf{r},\mathbf{v}_{s})-\mathbf{Y}_{c}(\mathbf{r})\|^{2}_{2}$
$+\alpha\underset{\mathbf{r}\in\mathcal{R}}{\sum}\|\mathbf{C}(\mathbf{r},\mathbf{v}_{s})-\hat{\mathbf{C}}(\mathbf{r},\mathbf{v}_{s})\|^{2}_{2}\cdot\mathbb{I}_{\mathbf{v}_{s}\notin\mathcal{T}^{m}},$ (4)
@@
$\mathbf{Y}_{c,t}(\mathbf{r}):=\sqrt{\alpha_{t}}\mathcal{E}_{v}(\mathbf{Y}_{c}(\mathbf{r}))+\sqrt{1-\alpha_{t}}\epsilon,$ (5)
$\hat{\mathbf{F}}_{s}(\mathbf{r},\mathbf{l}_{p})=\Theta_{u}(\mathbf{Y}_{c,t}(\mathbf{r}),\mathcal{E}_{c}(\mathbf{l}_{p})),$ (6)
where $\mathcal{E}_{v}$ is a VAE encoder to encode image $\mathbf{Y}_{c}$ from pixel space to latent semantic space. $t$ represents the diffusion process step, $\epsilon\sim\mathcal{N}(0,1)$ and $\alpha_{t}$ is designed to control the noise schedule. $\mathbf{l}_{p}$ denotes the language prompt modified from task description $\mathbf{l}$. To perform skill-shared semantic transfer, we align the rendered semantic feature $\mathbf{F}_{s}$ and diffusion feature $\hat{\mathbf{F}}_{s}$, and obtain the objective of our SSR module as follows:
$\mathcal{L}_{\mathrm{SSR}}=\mathcal{L}_{\mathrm{C}}+\lambda_{1}\mathcal{L}_{\mathrm{S}},$ (7)
$\mathcal{L}_{\mathrm{S}}=\underset{\mathbf{r}\in\mathcal{R}}{\sum}\|\mathbf{F}_{s}(\mathbf{r},\mathbf{v}_{s})-\hat{\mathbf{F}}_{s}(\mathbf{r},\mathbf{l}_{p})\|^{2}_{2},$ (8)
where $\lambda_{1}$ is the hyper-parameter.
III-C Skill-Shared Representation Distillation Module